Circular No. 8257

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 XD_{10} (LINEAR-NEAT) An apparently asteroidal object with not-unusual motion reported on

Dec. 4 and 5 by the LINEAR project, and designated 2003 XD_{10} on MPS 92917 (discovery observation below), was independently discovered with the NEAT 1.2-m Schmidt telescope at Palomar on Dec. 14.4 (discovery observation below) and reported then to be cometary (with a faint short tail toward the east-southeast) by K. J. Lawrence. Following posting on the NEO Confirmation Page, several other CCD observers have also noted the object’s cometary nature, including P. Birtwhistle (Great Shefford, Berkshire, England, 0.30-m reflector; on Dec. 14.9, from co-added images totaling 15 min exposure, diffuse coma of diameter 10′′, extended in p.a. ∼ 260^{o}, surrounding a central condensation of mag 19.4; on Dec. 15.9, 8′′ coma and 45′′ tail in p.a. 255^{o}), J. E. McGaha (Tucson, AZ, 0.36-m reflector; on Dec. 16.3, three co-added 1-min frames show a small starlike condensation with a 8′′ coma), J. Young (Table Mountain, CA, 0.6-m reflector; on Dec. 17.3, 3′′ asymmetric coma with a hint of tail ∼ 12′′ long in p.a. 250^{o}–260^{o}), and R. Fredrick and R. Trentman (Louisburg, KS, 0.75-m reflector; on Dec. 17.4, broad fan-shaped tail 20′′ long in p.a. 240^{o}).

2003 UT α_{2000} δ_{2000} Mag. Observer
Dec. 4.33554 6^{h}52^{m}35.^{s}22 +41^{o}38′56′′.6 19.2 LINEAR
14.43141 6 45 02.68 +43 35 49.0 20.1 NEAT

Additional astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2003-Y16.

T = 2003 Sept.16.8716 TT  ω = 9.^{o}1651
e = 0.436418  Ω = 43.5591 }2000.0
q = 1.881255 AU  i = 14.7293
a = 3.338032 AU n^{o} = 0.1616100 P = 6.1 years

SUPERNOVA 2003kz IN MCG +09-16-34 Further to IAUC 8245, T. Boles reports the discovery of an apparent

supernova (mag 18.5) on unfiltered CCD images taken on Dec. 16.092 and 17.186 UT. The new object is located at α = 9^{h}38^{m}35^{s}.27, δ = +52^{o}17′53′′.2, which is ≈ 9′′.5 west and 11′′.1 north of the center of MCG +09-16-34. SN 2003kz is not present on Boles’ images from Sept. 29 and Oct. 24 (limiting mag 19.5) or on Digitized Sky Survey red (1995) or blue (1993) plates.

2003 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green